A body starts from rest and travels with an acceleration of 2 m/s
1
Expert's answer
2021-07-19T00:27:01-0400

Gives

Initial velocity

u=0 m/sec2

Acceleration (a)=2m/sec2


Newton's motion 1st law

"v=u+at=0+2\times t"

"v=2t m\/sec"